Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2017-11-12

Observers: Francesca Onori, Giacomo Cannizzaro
TAT: Annalisa De Cia, Joe Lyman
Support Team: G. Cannizzaro, F. Onori, Z. Kostrzewa-Rutkowska, P. Jonker, K.M. Lopez
Instruments: SOFI, EFOSC2

Weather conditions:
UT 00:00 clear sky, humidity 45%, wind speed 2.2 m/s
UT 04:22 thin clouds, humidity 27%, wind speed 6.0 m/s, seeing 1.15

CALIBRATIONS:
EFOSC: bias, arcs (also for g20);


LOG
TARGET              TYPE                   UT                  FILTER/GRISM      SLIT          SEEING       AIRMASS       COMMENTS
Feige110              STD                       00:10               g11, g16 ,g13           1.0, 1.5                           1.1
AT2017iks            Follow up            00:40                BVgri                                            1.0               1.2
AT2017hxv          Classification      01:45                g13                            1.0              1.1               1.2
AT2017gww        Follow Up            01:58                g13                            1.0              1.2                                        Problem with focus
AT2017gww        Follow Up            02:07                g13                            1.0              1.0               1.3
AT2017hxz          Follow Up            02:47                g11                            1.0              1.0               1.2              
AT2017hxz          Follow Up            03:40                g16                            1.0               1.1              1.1
AT2017hyw          Classification      04:42                g13                            1.0               1.2              1.3                  

Switch to SOFI    UT 04:53
AT2017fzw          Follow Up           05:30                BG                              1.0               1.0              1.2                Difficult to center in the slit, object very faint in the field.
Hip31280            Telluric                07:00                BG, RG                       1.0               1.2              1.0
LTT3218              Standard             07:10                BG                              1.0               0.5              1.2               Allignment problems
LTT3218              Standard             07:21                BG                              1.0               0.6              1.2

Switch to EFOSC

ATLAS17naq      Classification      07:45               gr13                             1.0               0.5               1.2
L745a                  Standard             08:00             gr11,gr13,gr16             1.0,1.5        1.4               1.0
AT2017hyx         Classification      08:30              gr13                              1.0               0.5               1.3

Comments